# Exploring Knowledge Boundaries in Large Language Models for Retrieval Judgment

### v1

@@ 5.3 Use as a plug-in
We conduct an evaluation of several prominent LLMs using two benchmark datasets: WebQA and NaturalQA. For this analysis, we utilize the KBM as the retrieval judgment model and compare its performance against the LLM Only, RAG, and prompt methodologies applied to each LLM. Our findings reveal that KBM acts as a valuable plug-in, enhancing the end-to-end performance of other LLMs, although it falls short of achieving the comprehensive improvements offered by ALL RAG. This observation aligns with our previous analysis, which indicates that each LLM possesses a unique knowledge boundary, making it challenging to accurately represent the knowledge boundaries of all LLMs with a single model. The results are summarized in Table 5.2.
In particular, RAG demonstrates its capacity to enhance the response quality of LLMs, with a relatively stable upper limit for this improvement. While the effects on the LLM Only model vary significantly among the different LLMs, the scores converge when RAG is incorporated. The varying direct knowledge boundaries of LLMs contribute to the disparate enhancement effects observed with KBM. For instance, the KBM-Certainty judgment method shows that Qwen2-72B achieves an increase from 83.7% to 91.8%, representing an improvement of 8.1% points. In contrast, Llama3-70B exhibits a more substantial increase, improving from 48.7% to 88.2%, which corresponds to a remarkable gain of 39.5% points.
Overall, compared to the baseline, the knowledge boundary model typically enhances LLM performance across the test datasets. However, discrepancies exist between the known/unknown distributions of the knowledge boundary model and the general model. This divergence accounts for the varied performance enhancements observed among the different LLMs. For example, in robust LLMs, excessive unnecessary retrievals may occur. While Qwen2-72B may demonstrate a lower upper limit for retrieval, the knowledge boundary model still manages to execute 90% of the retrievals. Conversely, Llama3-70B encounters challenges in attaining the ALL RAG score even with a 90% retrieval rate in the WebQA dataset. Similar trends manifest in the English QA tests, suggesting that when the English proficiency of the KBM ontology model is lower than that of the generative LLM, performance improvements are more pronounced, and the reverse holds true as well.
5.4 Discussion
In this work, we find that while RAG effectively addresses certain challenges associated with LLMs, it is not universally applicable. As noted in Section §3.3, it can be counterproductive when relevant information is difficult to retrieve or noisy. This limitation is illustrated in Table 1, where RAG shows minimal performance improvement in tasks such as multiple-choice questions and math assessments.
We suggest that RAG’s effectiveness could be enhanced when deployed within specific modules through techniques such as question rewriting, page reading, and re-ranking. However, our primary focus is to investigate the knowledge boundaries of LLMs. Therefore, we consider strategies to mitigate the impact of RAG components that are neutral or detrimental, especially when LLM knowledge remains limited, thereby emphasizing RAG’s boundaries.
To achieve this objective, we implement the method illustrated in Figure 8, which aims to improve information processing while minimizing the influx of irrelevant input tokens and reducing noise from irrelevant retrieval results. Specifically, in specialized problem domains, such as multiple-choice questions, when the KBM identifies a query as unknown, we perform a search to retrieve concatenated chunks of relevant information. The KBM uses both the original query and the concatenated chunks as prompt input to assess whether the LLM can effectively answer the query with the additional retrieved information. If the LLM generates a response, the concatenated chunk is then forwarded to it; if not, the original query is transmitted directly. This method effectively minimizes the amount of information fed into the LLM (as indicated by the red area in Figure 8), while maintaining performance that is competitive with both LLM Only and ALL RAG scenarios, and in some instances, it even surpasses both approaches.
@@ 6 Conclusion
In this paper, we explore the impact of RAG on LLMs, depending on the types of questions they encounter. We find that a key limitation of existing approaches is that RAG often lacks adaptability to the knowledge boundaries of specific LLMs, which can lead to ineffective or even harmful responses. To address this challenge, we first classify the types of questions encountered by LLMs and identify appropriate search regions within the model boundaries. We then introduce a KBM for retrieval judgment, optimizing retrieval requests based on our proposed confidence and certainty methods. This model effectively classifies questions as known or unknown and determines whether retrieval is necessary. Our extensive evaluation on 11 English and Chinese datasets demonstrates that KBM significantly improves retrieval efficiency while maintaining high performance. It proves effective in three complex scenarios: dynamic knowledge, long-tail static knowledge, and multi-hop questions, showcasing its versatility in real-world applications. Our findings indicate that KBM, as an intelligent retrieval judgment strategy, effectively reduces resource consumption by decreasing retrieval calls, token usage, and response costs, while simultaneously enhancing the performance of LLMs.
Appendix A Appendix
Figure 9: en/zh- Prompt for Sampling LLM Only Response
Figure 10: en/zh- Prompt for Sampling RAG Response
Figure 11: en/zh- Prompt for Multiple Choice Question LLM Only Response
Figure 12: en/zh- Prompt for Multiple Choice Question RAG Response
Figure 13: zh-Prompt for Response Accuracy Assessment
Figure 14: KBM Prompt for English/Chinese Questions with/without RAG Demonstration
A.1 Implementation Details
@@
We employ four Nvidia A100 GPUs, each with 80GB of memory, to train our KBM models. Each model undergoes training for three epochs, utilizing a batch size of four. The peak learning rate is set to 1e-5, with a warmup ratio of 2% and cosine decay for the learning rate. To accommodate memory limitations, we restrict the maximum token length to 1580 for the 7B model and 1524 for the 13B model. For multi-GPU distributed training, we utilize Deepspeed stage 2 (Rajbhandari et al., 2020) while enabling Bfloat16 precision. Inference on the trained models is conducted using a single Nvidia Tesla V100 GPU with 32GB of memory.
A.2 Data Generation and KBM Inference Instructions
In this section, we show the instructions for producing data and KBM inference. Figure 9 shows the instructions for collecting LLM Only responses under QA tasks. Figure 10 shows the instructions for collecting LLM responses under RAG. Similarly, Figure 11 and Figure 12 show the instructions for collecting LLM Only/RAG in answering multiple-choice questions. It is also worth noting that since there are no abundant answers in the Chinese data QA data WebQA and SogouQA, directly using the full match method for evaluation will have a large number of errors, so we use Qwen2-72B and the evaluation prompt shown in Figure 13 for answer evaluation. In order to ensure the unbiasedness of the evaluation, we did not include the original question information in the instructions. During KBM training and inference, we adopt the prompt presented in Figure 14 and illustrate the dictionary reply format based on Mastery and Certainty.

### v2

@@ 4.3 Use as a Plug-in
To assess the effectiveness of the KBM model in triggering RAG, we apply it as a plug-in to GPT-4o, Llama3-70B, and Qwen2-72B. We evaluate these models using the WebQA and NQ test sets. The results are summarized in Table 4. Our findings indicate that KBM enhances the performance of these LLMs, although it does not achieve the comprehensive improvements provided by ALL RAG. This is consistent with our previous analysis showing that each LLM has a unique knowledge boundary, complicating the representation of knowledge boundary across all LLMs with a single model. The varying knowledge boundary of LLMs significantly influence the enhancement effects observed with knowledge boundary models (KBM). For example, the KBM-certainty judgment method shows that Qwen2-72B improves its performance from 83.7% to 91.8%, resulting in an increase of 8.1%, while Llama3-70B experiences a much larger gain, rising from 48.7% to 88.2%, equating to a 39.5% improvement. This illustrates that the initial knowledge capacity and boundary of an LLM can lead to divergent levels of enhancement, when interfacing with KBMs.
Additionally, RAG effectively enhances the response quality of LLMs, leading to a relatively stable upper limit for improvements across models. However, the performance gains differ due to the varying distributions of known and unknown knowledge in KBMs versus general models. For instance, although Llama3-70B demonstrates a notable retrieval rate of 90%, its upper limit is lower than that of Qwen2-72B, which struggles to reach the ALL RAG score despite the same retrieval performance in the WebQA dataset. These trends are echoed in English QA tests, where significant performance improvements are observed when the English proficiency of the KBM ontology model inferior to that of the generative LLM. This suggests that aligning the knowledge boundary with the capabilities of the LLM is essential for optimizing performance.
5 Related Work
Large Language Model Knowledge Exploration. The exploration of knowledge boundary in LLMs attracts significant attention. Kadavath et al. (2022) examines the self-evaluation capabilities of LLMs, showing that larger models enhance their calibration by initially proposing answers and then evaluating their validity. Ren et al. (2024) studies LLMs’ perception of factual knowledge boundary and finds that they often display blind confidence in their abilities. Yin et al. (2023) focuses on self-awareness, demonstrating that while LLMs can identify some unanswered questions, substantial discrepancies still exist, affecting their uncertainty detection. Chen et al. (2024) introduces COKE, an unsupervised method for teaching models to articulate their knowledge limits through internal signals, yielding improved outcomes across various datasets. Kang et al. (2024) points out that LLMs often default to examples in training data when facing unfamiliar queries. Li et al. (2024b) explores hallucinations related to insufficient prompt context, showing that models frequently fail to recognize inadequate information.
Retrieval-Augmented Generation. RAG enhances LLMs by integrating retrieved text passages, significantly improving performance in knowledge-intensive tasks. A key focus is optimizing the timing and strategy of retrieval. Asai et al. (2023) introduce SELF-RAG, a method that trains LLMs to retrieve information, generate content, and evaluate their outputs using reflection tokens. This method enables the customization of model behavior, demonstrating significant performance improvements over standard RAG approaches. Jeong et al. (2024) proposes Adaptive-RAG, which adjusts query handling based on complexity. Wu et al. (2024) explores how LLMs process erroneous retrieved content. By creating a dataset to assess model responses to incorrect information, the study reveals insights into how models correct their outputs or may perpetuate errors. Cuconasu et al. (2024) conducts a comprehensive study on the retriever’s function in RAG.
@@ 6 Conclusion
In this paper, we propose KBM to address the limitations of LLMs in managing dynamic knowledge and unknown static information. KBM judges whether to trigger RAG by indicating the known and unknown of a question, preserving RAG’s advantages while reducing computational costs. Validated on 11 datasets, KBM matches the performance of All RAG while reducing the retrieval ratio, enhancing efficiency in QA tasks without compromising other instruction fine-tuning performance. We also confirm KBM’s robustness in complex scenarios, such as dynamic and long-tail knowledge and multi-hop problems, highlighting its effectiveness as an external LLM plug-in. Our contributions include methods for generating soft labels based on accuracy and certainty, enabling LLMs to better express their knowledge status.
Limitations
Our work is the first attempt to explore how to reduce the proportion of retrievals that trigger RAG while maintaining high performance. Despite the reduction in the proportion of triggers, our study has some limitations that may guide future efforts.
First, we propose a method based on accuracy and certainty sampling to generate soft labels for training KBM, demonstrating that LLMs learn to express unknowns in different tasks without requiring a large amount of data. However, in real-world scenarios, it is essential to include more diverse QA format data for better generalization, such as reasoning and content generation. Additionally, converting data from various question-answer formats into the format of QA pairs enhances the performance of KBM.
Secondly, while we enable LLMs to express unknowns and trigger RAG, there remains a relatively low probability that an unknown question posed by LLM can be resolved by RAG, particularly when considering post-retrieval operations such as changing the retriever, sorting, and denoising.
Appendix A Additional Experimental Analysis
A.1 Analysis of KBM in Sub-intervals
We divided the accuracy interval into steps of 0.2 on the Chinese WebQA test set to investigate the improvements of KBM compared to RAG. As illustrated in Figure 8, our findings indicate that the KBM method, based on Accuracy and Certainty, demonstrates significant enhancements across several intervals, particularly in the lower accuracy ranges. For instance, when the LLM-only accuracy is at 0.0, KBM-Certainty shows an increase of 79.4%, while KBM-Accuracy exhibits a rise of 74.2%. Notably, when compared to All RAG, KBM achieves performance levels that are comparable to RAG in multiple intervals, especially within the Certainty-based approach.
A.2 Case Study
As illustrated in Figure 7, we showcase how KBM functions within a QA context. For questions (1) and (4), if KBM determines that the answer is Known, the LLM can provide the correct response directly, eliminating the need to search for external resources. In the case of question (2), which the LLM cannot answer or chooses not to address, KBM can preemptively classify it as unknown. This classification enables subsequent actions to penalize the search tool, prompting it to retrieve relevant textual materials that can then be provided to the LLM as context to aid in formulating an answer. For questions (3) and (5), direct use of the LLM to respond may lead to hallucinations, resulting in the output of incorrect information. In these instances, KBM can classify the question as unknown and activate RAG to produce a more reliable answer.
Search API Token time
A/E Call Rate consumption (ms)
All RAG 79.16 100.00% 593 2498
KBM-Acc. 78.77 67.50% 421 2231
KBM-Cer. 79.22 86.50% 517 2384
Table 5: Efficiency analysis across 11 datasets.
A.3 Efficiency analysis
We evaluate our method across 11 datasets and compare it to the All RAG approach. The findings appear in Table 5. For KBM-Acc, the API calls reduce by 32.5%; tokens input to LLM decrease from 593 to 421; and the total computation time reduces from 2498 ms to 2231 ms. These results indicate that KBM-Acc offers significant advantages in both computational and invocation costs compared to All RAG. On the other hand, when comparing with All RAG, our KBM-Cer reduces all three metrics, namely API calls, token consumption, and time, to varying degrees. Importantly, KBM-Cer not only controls costs but also surpasses All RAG in overall performance metrics.
Appendix B Implementation Details
B.1 Parameter Settings
@@
We employ four Nvidia A100 GPUs, each with 80GB of memory, to train our KBM models. To maintain high end-to-end performance, we set the threshold for both Accuracy and Certainty at 0.9. Each model undergoes training for three epochs, utilizing a batch size of four. The peak learning rate is set to 1e-5, with a warmup ratio of 2% and cosine decay for the learning rate. To accommodate memory limitations, we restrict the maximum token length to 1580 for the 7B model and 1524 for the 13B model. For multi-GPU distributed training, we utilize Deepspeed stage 2 (Rajbhandari et al., 2020) while enabling Bfloat16 precision. Inference on the trained models is conducted using a single Nvidia Tesla V100 GPU with 32GB of memory.
B.2 Datasets
English Datasets. NaturalQA (Kwiatkowski et al., 2019): This QA dataset, curated by Google, consists of real-world questions derived from natural retrieval queries. TriviaQA (Joshi et al., 2017): This dataset is based on encyclopedic content and features complex questions and answers, primarily sourced from competitions and quizzes. MMLU (Hendrycks et al., 2021): This dataset comprises multiple-choice questions across various fields, assessing the model’s knowledge mastery in academic and professional domains. OpenBookQA (Mihaylov et al., 2018): Focusing on scientific inquiries, this dataset requires reasoning rooted in principles and common sense. en-SQuAD-en2.0 (Rajpurkar et al., 2018): This dataset features question-answer pairs and evaluates reading comprehension skills. FreshQA-en (Vu et al., 2023): This dataset presents various question and answer types, offering a comprehensive assessment of QA capabilities. HotpotQA (Yang et al., 2018): This dataset consists of 113,000 Wikipedia based QA pairs that necessitate complex reasoning across multiple supporting documents and include sentence-level supporting facts.
Chinese Datasets. For the Chinese dataset, we utilize the following resources: WebQA (Chang et al., 2022): This open-domain QA dataset is collected via web crawlers, covering a wide array of topics and evaluating the performance of QA systems. SogouQA  [^5] https://github.com/sherlcok314159/ChineseMRC-Data: Provided by Sogou, this dataset features user-generated questions and system-generated answers, assessing accuracy and robustness. MLEC (Li et al., 2021): This dataset is designed to test the comprehension capabilities of models in various contexts. Xiezhi (Gu et al., 2024): A set of 249,587 Chinese/English questions covering 516 subjects for evaluating LLMs. SQuAD-zh  [^6] https://github.com/pluto-junzeng/ChineseSquad: The Chinese version of the English SQuAD (Rajpurkar et al., 2018) dataset serves to train and evaluate machine reading comprehension and QA systems. Chinese SimpleQA(CSQA) (He et al., 2024): A benchmark for evaluating the factuality of LLMs in short Chinese QA across diverse topics. CMMLU (Li et al., 2024a):This comprehensive benchmark assesses the knowledge and reasoning capabilities of language models across 67 topics, from basic to advanced.